# A Multi-Stream Fusion Approach with One-Class Learning for Audio-Visual Deepfake Detection

### v1

@@ V-D Impact of Visual Feature Extractor
Table II demonstrates that models with Scnet-STIL visual feature extractor perform better on the RAFV test set. Thus, this section examines the impact of visual feature extractors in one-class learning. Although OC-Softmax effectively compacts genuine representations and distributes fake representations, its performance is limited if the visual feature extractor fails to capture the general features of fake visual artifacts. This limitation arises because OC-Softmax compacts real representations based on observed attacks and real data, potentially including unseen attack representations within the realm of genuine representations. Therefore, extracting more general features of fake videos, such as Spatial-Temporal Inconsistency, could be beneficial. Fig. 4 compares the visual scores from both visual feature extractors. We can observe that the ResNet-based visual feature extractor lacks the ability to detect unseen fake methods effectively compared to the SCNet-STIL-based visual feature extractor. This explains why models with the STIL feature extractor significantly outperform models with a ResNet feature extractor on the RAFV test set.
@@ VI Conclusion
This paper presents a multi-stream fusion framework with one-class learning to enhance audio-visual deepfake detection. Our proposed framework improves detection performance against unseen deepfake generation methods by an average of 7.31% over the baseline model across four test sets. Additionally, the MSOC framework provides interpretability, offering the ability to identify which modality is fake, which can be achieved through the score distribution of the models (Audio, Visual, Audio-Visual). Future work includes joint modeling of detecting audio-visual unsynchronization and deepfakes and a more robust framework for rooting the fake modality.

### v2

@@ V-D Impact of Visual Feature Extractor
Table II demonstrates that models with SCNet-STIL visual feature extractor perform better on the RAFV test set. Thus, this section examines the impact of visual feature extractors in one-class learning. Although OC-Softmax effectively compacts genuine representations and distributes fake representations, its performance is limited if the visual feature extractor fails to capture the general features of fake visual artifacts. This limitation arises because OC-Softmax compacts real representations based on observed attacks and real data, potentially including unseen attack representations within the realm of genuine representations. Therefore, extracting more general features of fake videos, such as Spatial-Temporal Inconsistency, could be beneficial.
Fig. 4 compares the visual scores from both visual feature extractors. We can observe that the ResNet-based visual feature extractor lacks the ability to detect unseen fake methods effectively compared to the SCNet-STIL-based visual feature extractor. This explains why models with the STIL feature extractor significantly outperform models with a ResNet feature extractor on the RAFV test set.
@@ VI Conclusion
This paper presents a multi-stream fusion framework with one-class learning to enhance audio-visual deepfake detection. Our proposed framework improves detection performance against unseen deepfake generation methods compared to SOTA models. Additionally, the MSOC framework provides interpretability, offering the ability to identify which modality is fake, which can be achieved through the score distribution of the models (Audio, Visual, Audio-Visual). Future work includes joint modeling of detecting audio-visual unsynchronization and deepfakes and a more robust framework for rooting the fake modality.